OpenStudy (anonymous):

Can someone help me find x&y for these equations? x-3y=6 3x=6y+18

OpenStudy (australopithecus):

x = 6 + 3y 3x = 6y + 18 3(6+3y) = 6y + 18 solve for y then sub y into one of the original equations

OpenStudy (anonymous):

Okay, thanks. :)